# PROV-AGENT: Unified Provenance for Tracking

### v2

@@ Abstract
Large Language Models (LLMs) and other foundation models are increasingly used as the core of AI agents. In agentic workflows, these agents plan tasks, interact with humans and peers, and influence scientific outcomes across federated and heterogeneous workflows. However, agents can hallucinate or reason incorrectly, propagating errors when one agent’s output becomes another’s input. Thus, assuring that agents’ actions are transparent, traceable, reproducible, and reliable is critical to assess hallucination risks and mitigate their workflow impacts. While provenance techniques have long supported these principles, existing methods fail to capture and relate agent-centric metadata such as prompts, responses, and decisions with the broader workflow context and downstream outcomes. In this paper, we introduce PROV-AGENT, a provenance model that extends W3C PROV and leverages the Model Context Protocol (MCP) and data observability to integrate agent interactions into end-to-end workflow provenance. Our contributions include: (1) a provenance model tailored for agentic workflows, (2) a near real-time, open-source system for capturing agentic provenance, and (3) a cross-facility evaluation spanning edge, cloud, and HPC environments, demonstrating support for critical provenance queries and agent reliability analysis.
@@
Workflows, Agentic Workflows, Provenance, Lineage, Responsible AI, LLM
@@ V Conclusion
Acknowledgments. The authors would like to thank the ORNL team: Miaosen Chai, Timothy Poteet, Phillipe Austria, Marshall McDonnell, Ross Miller, A.J. Ruckman, Tyler Skluzacek, Feiyi Wang, Sarp Oral, Arjun Shankar for their assistance and guidance with the use case development. ChatGPT-4o was used to help polish writing, improve conciseness, and check grammar across the sections of the paper. This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725.

### v3

@@ Abstract
Large Language Models (LLMs) and other foundation models are increasingly used as the core of AI agents. In agentic workflows, these agents plan tasks, interact with humans and peers, and influence scientific outcomes across federated and heterogeneous environments. However, agents can hallucinate or reason incorrectly, propagating errors when one agent’s output becomes another’s input. Thus, assuring that agents’ actions are transparent, traceable, reproducible, and reliable is critical to assess hallucination risks and mitigate their workflow impacts. While provenance techniques have long supported these principles, existing methods fail to capture and relate agent-centric metadata such as prompts, responses, and decisions with the broader workflow context and downstream outcomes. In this paper, we introduce PROV-AGENT, a provenance model that extends W3C PROV and leverages the Model Context Protocol (MCP) and data observability to integrate agent interactions into end-to-end workflow provenance. Our contributions include: (1) a provenance model tailored for agentic workflows, (2) a near real-time, open-source system for capturing agentic provenance, and (3) a cross-facility evaluation spanning edge, cloud, and HPC environments, demonstrating support for critical provenance queries and agent reliability analysis.
@@
Workflows, Agentic Workflows, Provenance, Lineage, Responsible AI, LLM, Agentic AI
@@ V Conclusion
Acknowledgments. The authors thank the ORNL team: Miaosen Chai, Timothy Poteet, Phillipe Austria, Marshall McDonnell, Ross Miller, A.J. Ruckman, Tyler Skluzacek, Feiyi Wang, Sarp Oral, Arjun Shankar for their help with the use case development. ChatGPT-4o was used to help polish writing, improve conciseness, and check grammar. This research used resources of the Oak Ridge Leadership Computing Facility at ORNL, supported by the Office of Science, U.S. Department of Energy (DOE) under Contract No. DE-AC05-00OR22725. Additional DOE support was provided under Contract No. DE-AC02-06CH11357.